Circular No. 8574

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA SCORPII 2005 G. Pojmanski, Warsaw University Astronomical Observatory, reports

the ASAS discovery (cf. IAUC 8402, 8495) of a possible nova on CCD images obtained with a 180-mm-focal-length (65-mm-aperture) telephoto lens, giving the new object’s position as α = 17^{h}44^{m}22^{s}, δ = −34^{o}16′.5 (equinox 2000.0 presumed), which is noted as being in an area crowded with nearby stars; V magnitudes for the new object: July 23.287 UT, [14; 25.284, 9.11; 26.296, 8.98; 26.311, 8.98. S. Nakano, Sumoto, Japan, reports the independent discovery of this new object at mag 8.7 by Hideo Nishimura (Kakegawa, Shizuoka-ken, Japan) on two 30-s Kodak T-Max films taken on July 26.565 with a Pentax 6×7 camera (+ 200-mm f/4.0 lens), with position end figures given as 22^{s}, 15′59′′ (equinox 2000.0); nothing was visible at this location on Nishimura’s survey films taken on July 20 (limiting mag 11). Nakano also forwards the following precise position from four unfiltered 20-s CCD frames taken by N. Hashimoto and T. Urata, Bisei Spaceguard Center, around July 26.596 using the 0.50-m f/2.0 reflector: α = 17^{h}44^{m}21^{s}.59, δ = −34^{o}16′35′′.7 (magnitude 8.6). Nothing is visible at this location on Digitized Sky Survey images (no limiting magnitude provided); the nearest USNO-A2.0 star (red mag 17.4) has position end figures 21^{s}.987, 35′′.85.

VARIABLE STARS NEAR GALAXIES Further to IAUC 8570, M. Baek and W. Li report the LOSS discovery

of a variable star on unfiltered KAIT images taken on July 22.35 (at mag 15.6), 23.35 (mag 15.6), and 24.40 UT (mag 15.7). The variable star is located at α = 21^{h}43^{m}32^{s}.94, δ = +46^{o}12′22′′.2 (equinox 2000.0), which is 48′′.8 east and 108′′.9 south of the nucleus of UGC 11800. A KAIT image taken on July 16.36 showed nothing at this position (limiting mag ∼ 19.0).

E. Lee and W. Li also report the LOSS discovery of a variable star located 111′′.1 east and 9′′.7 north of the nucleus of MCG –04-35-5 on unfiltered KAIT images taken on July 25.19 (when the new object was at mag 15.8) and 26.18 UT (mag 15.9). This second variable is located at α = 14^{h}43^{m}35^{s}.30, δ = −23^{o}26′30′′.6 (equinox 2000.0). A KAIT image taken on 2004 May 24.20 showed nothing at this position (limiting mag ∼ 19.0).

COMET 168P/2005 N2 (HERGENROTHER) This comet (P/2005 N2 = 1998 W2; cf. IAUC 8560) has been given

the permanent numbering 168P (MPC 54523).

2005 July 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green